find a polynomial equation with integer coefficients for the set of coplanar points described;
each point is twice as far from the line x=3 as it is from the line y =-2
also;
is the graph a conic section, if so what section?

Answer by stanbon(75887)

find a polynomial equation with integer coefficients for the set of coplanar points described;
each point is twice as far from the line x=3
as it is from the line y =-2
------
also;
is the graph a conic section, if so what section?
--
Let the points on x = 3 be (3,y)
---------------------------------------
Let the points on y = -2 be (x,-2)
-----
Equation:
(x,y)'s distance from (3,y) = 2*(x,y)'s distance from (x,-2)
-----
sqrt[(x-3)^2 + 0] = 2*sqrt[0 + (y-2)^2]
(x-3)^2 = 4(y-2)^2
-----
(x-3)^2/4 - (y-2)^2 = 0
-----
etc.
Cheers,
Stan H.